Be sure to answer all parts How many total moles of ions are released when the following sample dissolves completely in water? Enter vour answer in scientific notation. 9.75 x 10-5 g of Ca(NO3)2: ×10 mol
Answer #1

Molar mass of Ca(NO3)2,

MM = 1*MM(Ca) + 2*MM(N) + 6*MM(O)

= 1*40.08 + 2*14.01 + 6*16.0

= 164.1 g/mol

mass(Ca(NO3)2)= 9.75*10^-5 g

number of mol of Ca(NO3)2,

n = mass of Ca(NO3)2/molar mass of Ca(NO3)2

=(9.75*10^-5 g)/(164.1 g/mol)

= 5.941*10^-7 mol

1 molecule of Ca(NO3)2 has 3 ions

So,

number of mol ions = 3*5.941*10^-7 mol

= 1.78*10^-6 mol

Answer: 1.78*10^-6 mol